# Investigating cannibalistic millisecond pulsar binaries using MESA : new constraints from pulsar spin and mass evolution

### v1

@@ 3.2 Orbital evolution and surface hydrogen abundance
There is another source in the same observed parameter space as the two tidarrens (see Figure 3), for which the surface H abundance is unknown: PSR J0636+5128 (Stovall et al., 2014; Draghis & Romani, 2018). Since it is in the same region of the parameter space as PSR J1311–3430 and PSR J1653–0159 and lies on similar evolutionary tracks, it is very likely this source had a similar formation history and will be H deficient on its surface. Since tidarrens can be produced with any value of $f_{\rm pulsar}$, the effect of irradiation on their evolution is not significant. They reach their extremely low masses due to the extended RLO phase along with stellar winds during the giant phase, during which they lose their entire H envelope. As we saw in Figure 3, the lower-right part of the RB population is also reproduced. These systems are the very same that pass through the UCXB region and lead to the tidarrens, and correspond to systems that start their RLO just below $P_{\rm bif}$. As there is no irradiation included in their evolution, these systems could depict some of the non-irradiated RBs observed. However, since all currently observed BWs and about half of the currently observed RBs are strongly irradiated (Turchetta et al., 2023), we investigate higher values of $f_{\rm pulsar}$ to form RBs in the subsequent sections.
@@
Figure 4: Allowed initial parameter space to form spider MSPs, for an initial NS of mass $1.3\,{\rm M}_{\odot}$ and $f_{\rm pulsar}$ values of 0.0 (left) and 0.5 (right). $M^{i}_{\rm c}$ is the initial mass of the companion to the NS and $P^{i}_{\rm orb}$ is the initial orbital period. The colorbar shows the final orbital periods ($P^{f}_{\rm orb}$) and the gray symbols did not attain MSP-like spins. We show the end states of the tracks, if they encountered a dynamical instability (triangle symbols) or ended with one of the criteria for successful termination (star symbols). The diverging symbols are shown by gray crosses. We also show the near-$P_{\rm bif}$ binaries that passed through the RB part of the observed parameter space enclosed by dashed orange lines. The magenta triangle shows the initial parameters of the binary that reproduces PSR J0952–0607 (with observed spin period of 1.41 ms).
@@
Figure 5: Orbital period evolution versus companion mass of selected binaries from our simulated grid shown in Figure 1. $M_{\rm c}$ is the mass of the companion to the NS and $P_{\rm orb}$ is the orbital period. The figure shows the effect of pulsar wind irradiation, comparing $f_{\rm pulsar}=0.0$ (dashed lines) and 0.5 (solid lines). Left panel shows the tracks with $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$, $P^{i}_{\rm orb}=1.0$, 2.4, and 2.6 d. Right panel shows two initial companion masses, $M^{i}_{\rm c}=1.0$ (with $P^{i}_{\rm orb}=1$ d) and $2.5\,{\rm M}_{\odot}$ ($P^{i}_{\rm orb}=2.0$ and 2.2 d). We compare our tracks to $M_{\rm c,min}$ and $P_{\rm orb}$ for BWs (faint black crosses), RBs (faint red crosses), and huntsmen spiders (faint green crosses), from Nedreaas (2024, and references therein). The green star is the huntsman candidate 2FGL J0846.0+2820, observed by Fermi-LAT (Swihart et al., 2017). The magenta stars show the three transitional MSPs (tMSPs) observed, PSR J102347.67+003841.2 ($M_{\rm c}=0.22\,{\rm M}_{\odot}$ and $P_{\rm orb}=4.75$ hr; Shahbaz et al., 2019; Thorstensen & Armstrong, 2005), IGR J18245–2452 ($M_{\rm c,min}=0.17\,{\rm M}_{\odot}$ and $P_{\rm orb}=11.03$ hr; Papitto et al., 2013), and XSS J12270-4859 ($M_{\rm c,min}=0.15\,{\rm M}_{\odot}$ and $P_{\rm orb}=6.91$ hr; Roy et al., 2015; de Martino et al., 2015; Bassa et al., 2014).
@@
Figure 6: Orbital period versus companion mass for diverging binary tracks from our simulated grids that have similar NS spin as the observed PSR J0952–0607 (1.41 ms) within a 10 % margin (dashed gray curves). $M_{\rm c}$ is the mass of the companion to the NS and $P_{\rm orb}$ is the orbital period. All the tracks have $\beta=0.3$ and $f_{\rm pulsar}=0.5$. We further show the binary that best matches the sources in its orbital evolution in blue curve, which has initial companion mass of 2.0 M, initial NS mass of 1.3 M, and initial orbital period of 1.0 d.
@@ 3.3 Effect of pulsar wind irradiation
Figure 1 shows the evolutionary grids with an $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$ and varying pulsar wind efficiency ($f_{\rm pulsar}$). We also show the initial parameter space for spiders in Figure 4, their dynamical end, as well as the final orbital periods in the color bar, for $f_{\rm pulsar}=0.0$ and 0.5 and initial NS masses $1.3\,{\rm M}_{\odot}$. Many of the binaries shown in Figure 4 (around 50% in simulated spiders) end up suffering a dynamical instability (shown by diamond symbols in the figure); the companion exceeds its Roche lobe to fill out the gravitational equipotential surface passing through the second Lagrange point (Misra et al., 2020). This is a dynamically unstable process and leads to the binary merging.
@@
The companion at the end of the first RLO phase is a fully convective star. Convective envelopes are known to expand rapidly on mass loss (Hjellming & Webbink, 1987), driving sudden high mass-transfer rates during the second RLO phase and a dynamical instability. This occurs with and without pulsar wind irradiation. Irradiation removes some matter from the surface of the companion star and causes rapid stellar expansion. Increased $f_{\rm pulsar}$ results in a higher number of dynamical unstable systems (goes from 50% to 62% in simulated spiders when going from 0.0 to 0.5 in $f_{\rm pulsar}$). Despite this, we include these dynamically unstable systems in our study since by the time the instability onset happens, the average MSP binary is already $\sim 10^{1}0\rm\,yrs$ old and could have been observed during the preceding MSP phase. Hence, there is a reasonable chance to observe these binaries as BW binaries within a Hubble time ($1.4\times 10^{10}\rm\,yrs$). For a part of the parameter space the companion star would be destroyed due the instability. Therefore, it could provide an additional channel to the population of Galactic isolated MSPs (20% of the total Galactic MSP population) that is observed (Manchester et al., 2005a), with previous RLO spinning up the NS and subsequent merging destroying the companion.
@@
Going from $f_{\rm pulsar}=0.0$ to $f_{\rm pulsar}=0.5$ in the converging systems (Figure 4), the final orbital periods widen significantly and this effect depends on the initial properties of binary. The widest orbits are seen with $M^{i}_{\rm c}=1.0\,{\rm M}_{\odot}$, where the final orbital periods are $\gtrsim 10$ d. For higher companion masses, the final orbital periods lie in the range of 0.1 to 1 d. Figure 5 shows the orbital evolution corresponding to three $M^{i}_{\rm c}$ (1.0, 1.5, and 2.5 M) and $f_{\rm pulsar}=0.0$ to $f_{\rm pulsar}=0.5$, for various values of $P^{i}_{\rm orb}$. For case A RLO binaries ($P^{i}_{\rm orb}=1.0$ d; shown by black lines in the left panel), we see increased orbital widening with $f_{\rm pulsar}=0.5$. The companions to the NS become fully convective around 0.2 M, after which the pulsar wind is turned on. The orbit responds to the mass loss by expanding depending on $f_{\rm pulsar}$. Going from $M^{i}_{\rm c}=1.0\,{\rm M}_{\odot}$ to $1.5\,{\rm M}_{\odot}$ (for $f_{\rm pulsar}=0.5$) in right and left panels, respectively, the expanding orbits cover a larger part of the observed BW space. This is due the fact that all of these binaries are undergoing RLO on a nuclear timescale. For a fixed NS mass and orbital period, more massive companions (or higher $q$) lead to more accretion onto the NS and a higher spin-down luminosity (and correspondingly high orbital expansion). However, with an $M^{i}_{c}=4.0\,{\rm M}_{\odot}$, the accreted matter decreases down to about $0.004\,{\rm M}_{\odot}$ as the binary suffers a dynamical instability soon after the first RLO-onset. This is significantly less than $1.0\,{\rm M}_{\odot}$ accreted with $M^{i}_{c}=3.0\,{\rm M}_{\odot}$. Hence, there is a limit to how high in companion mass we can go with a fixed initial NS mass and reproduce BWs. We can see this clearly by looking at Figure 4.
@@
Next we look at the tracks for $P^{i}_{\rm orb}=2.4$ d in Figure 5, corresponding to the gray curves (left panel). For $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$, the effect of $f_{\rm pulsar}=0.5$ is significant on the orbital evolution. The near-$P_{\rm bif}$ binary maintains a wide orbit (around 0.2 d) crossing the RB regime. Going to $M^{i}_{\rm c}=2.5\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=2.0\,d$ (gray curves in right panel), the near-$P_{\rm bif}$ tracks no longer pass through the RB region with the heavier companion (irrespective of $f_{\rm pulsar}$). For intermediate-mass companions ($\gtrsim 2.0\,{\rm M}_{\odot}$), the initial angular momentum loss by RLO dominates over magnetic braking. For very short durations, the mass-transfer rate could even breach the Eddington limit, which results in any excess material being lost from the system taking away even more angular momentum. This continues until the star reaches around $1.0\,{\rm M}_{\odot}$ (and 1 d in Figure 5) and magnetic braking starts to dominate. The following evolution is governed by magnetic braking losses and behaves similar to case A RLO, during its first RLO phase. The fast initial mass loss causes the initially $2.5\,{\rm M}_{\odot}$ star to become fully convective around $0.2\,{\rm M}_{\odot}$, after which it follows a UCXB-like evolution. Hence, initial companion masses $\gtrsim 2.0\,{\rm M}_{\odot}$ in near-$P_{\rm bif}$ binaries do not reproduce RBs (also shown in Figure 4). If a heavy NS is present ($\sim 2.0\,{\rm M}_{\odot}$), it could be argued that correspondingly heavier companions ($\gtrsim 2.0\,{\rm M}_{\odot}$) could form RBs. However, NSs with masses $1.8\,{\rm M}_{\odot}$ and above very easily cross the maximum NS mass threshold and very few binaries in those grids would be viable spiders (see Section 3.1).
@@
Figure 7: Distribution of the final NS spins ($P^{f}_{\rm spin}$) in diverging binaries, for $\beta=0.0$ (top), $\beta=0.3$ (middle), and $\beta=0.7$ (bottom). We group the binaries by the initial NS mass ($M^{i}_{\rm NS}$) and also compare the spins to the observed spin of the huntsmen source 1FGL J1417.7–440 (2.66 ms, shown by the vertical gray dashed lines; Camilo et al., 2016).
Figure 8: Distribution of the final NS masses ($M^{f}_{\rm NS}$) in diverging binaries (for $\beta=0.0$), grouping the binaries by the initial NS mass ($M^{i}_{\rm NS}$). We also compare to the estimated pulsar masses of the two huntsmen sources 1FGL J1417.7–440 ($1.62^{+0.43}_{-0.17}\,{\rm M}_{\odot}$, shown by the vertical gray dashed lines and gray shaded area for the uncertainties; Swihart et al., 2018) and 2FGL J0846.0+2820 ($1.96^{+0.41}_{-0.41}\,{\rm M}_{\odot}$, shown by the vertical blue dashed lines and blue shaded area for the uncertainties; Swihart et al., 2017).
Figure 9: Allowed initial parameter space to form huntsmen spiders ($f_{\rm pulsar}=0.0)$, for an initial NS of mass $1.3\,{\rm M}_{\odot}$ (left) and $2.0\,{\rm M}_{\odot}$ (right). $M^{i}_{\rm c}$ is the initial mass of the companion to the NS and $P^{i}_{\rm orb}$ is the initial orbital period. The colorbar shows the final NS spins ($P^{f}_{\rm spin}$), the gray stars denoting systems that formed diverging orbits but did not spin-up to the MSP range. We show the end states of the tracks, if they encountered a dynamical instability (triangle symbols) or ended with one of the criteria for successful termination (star symbols). Converging tracks are shown by gray crosses.
@@
Table 1: Observed and estimated properties of the two observed huntsmen spiders. $M_{\rm c,min}$ is the minimum mass, calculated assuming a pulsar mass of $1.4\,{\rm M}_{\odot}$ and inclination of $90\degree$, $P_{\mathrm{orb}}$ is the orbital period, $P_{\rm spin}$ is the observed pulsar spin, $M_{\mathrm{NS}}$ is the estimated pulsar mass, $M_{\mathrm{c}}$ is the companion mass observed from Fermi-LAT, and $e$ is the observed eccentricity.

### v2

@@ 3.2 Orbital evolution and surface hydrogen abundance
The binary that reaches the minimum $P_{\rm orb}$ (19 min; Figure 3 top left panel) in the grid slice (with an $M^{i}_{\rm c}=1.0\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=2.6$ d) forms a fully convective companion after the companion reaches $\lesssim 0.02\,{\rm M}_{\odot}$. The NS present for this binary also accretes the least in our selected binaries, and the NS spin-up occurs the latest. Toward the end of the second RLO phase, while the companion star is detaching from its Roche lobe, orbital angular momentum losses due to gravitational radiation are still operating (strong for periods $\lesssim 1\,$d). This causes the companion star to fill its Roche lobe again. As the companion is fully convective, it responds by expanding (Hjellming & Webbink, 1987) further extracting angular momentum from the orbit (due to spin-orbit coupling). As both angular momentum losses (due to RLO and spin-orbit coupling) reduce the binary separation, the Roche-lobe radius is proportionally affected and the companion overfills its Roche lobe, increasing the mass-transfer rate to $\lesssim 10^{-2}\,{\rm M}_{\odot}\rm\,yr^{-1}$ and triggering a dynamical instability.
@@
There is another source in the same observed parameter space as the two tidarrens (see Figure 1), for which the surface H abundance is unknown: PSR J0636+5128 (Stovall et al., 2014; Draghis & Romani, 2018). Since it is in the same region of the parameter space as PSR J1311–3430 and PSR J1653–0159 and lies on similar evolutionary tracks, it is very likely this source had a similar formation history and will be H deficient on its surface. Since tidarrens can be produced with any value of $f_{\rm pulsar}$, the effect of irradiation on their evolution is not significant. They reach their extremely low masses due to the extended RLO phase along with stellar winds during the giant phase, during which they lose their entire H envelope. As we saw in Figure 3, the lower-right part of the RB population is also reproduced. These systems are the very same that pass through the UCXB region and lead to the tidarrens, and correspond to systems that start their RLO just below $P_{\rm bif}$. As there is no irradiation included in their evolution, these systems could depict some of the non-irradiated RBs observed. Since all currently observed BWs and about half of the currently observed RBs are strongly irradiated (Turchetta et al., 2023), we investigate higher values of $f_{\rm pulsar}$ to form RBs in the subsequent sections.
@@
Figure 4: Allowed initial parameter space to form spider MSPs for an initial NS of mass $1.3\,{\rm M}_{\odot}$ and $f_{\rm pulsar}$ values of 0.0 (left) and 0.5 (right). The term $M^{i}_{\rm c}$ is the initial mass of the companion to the NS, and $P^{i}_{\rm orb}$ is the initial orbital period. The color bar shows the final orbital periods ($P^{f}_{\rm orb}$) and the gray symbols did not attain MSP-like spins. We show the end states of the tracks, if they encountered a dynamical instability (with gray triangles or with colorful triangles depending on if onset of instability occurred during the first RLO phase or the second, respectively) or ended with one of the criteria for successful termination (star symbols). The diverging symbols are shown by gray crosses. We also show the near-$P_{\rm bif}$ binaries that passed through the RB part of the observed parameter space enclosed by dashed orange lines. The magenta triangle shows the initial parameters of the binary that reproduces PSR J0952–0607 (with observed spin period of 1.41 ms).
Figure 5: Evolution of the mass-loss rates ($\dot{M}_{\rm c}$, black curves) and the mass-transfer rates ($\dot{M}_{\rm transfer}$, blue curves) of the companion star. The initial binary masses are $M^{i}_{\rm c}=1.0$${\rm M}_{\odot}$and $M^{i}_{\rm NS}=1.3$${\rm M}_{\odot}$, with $\beta=0.3$. We present the three cases of interacting binaries studied, case A (top row; with $P^{i}_{\rm orb}=1.0$ d), near-$P_{\rm bif}$ (middle row; with $P^{i}_{\rm orb}=2.4$ d), and case B (bottom row; with $P^{i}_{\rm orb}=2.6$ d), each for three values of $f_{\rm pulsar}=0.05$ (left column), 0.1 (middle column), and 0.5 (right column).
@@
Figure 6: Orbital period evolution versus companion mass of selected binaries from our simulated grid shown in Figure 1. The term $M_{\rm c}$ is the mass of the companion to the NS, and $P_{\rm orb}$ is the orbital period. The figure shows the effect of pulsar wind irradiation, comparing $f_{\rm pulsar}=0.0$ (dashed lines) and 0.5 (solid lines). The left panel shows the tracks with $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$, $P^{i}_{\rm orb}=1.0$, 2.4, and 2.6 d. The right panel shows two initial companion masses, $M^{i}_{\rm c}=1.0$ (with $P^{i}_{\rm orb}=1$ d) and $2.5\,{\rm M}_{\odot}$ ($P^{i}_{\rm orb}=2.0$ and 2.2 d). We compared our tracks to $M_{\rm c,min}$ and $P_{\rm orb}$ for BWs (faint black crosses), RBs (faint red crosses), and huntsmen spiders (faint green crosses), from Nedreaas (2024, and references therein). The green star is the huntsman candidate 2FGL J0846.0+2820, observed by Fermi-LAT (Swihart et al., 2017). The magenta stars show the three transitional MSPs (tMSPs) observed, PSR J1023+0038 ($M_{\rm c}=0.22\,{\rm M}_{\odot}$ and $P_{\rm orb}=4.75$ hr; Shahbaz et al., 2019; Thorstensen & Armstrong, 2005), IGR J18245–2452 ($M_{\rm c,min}=0.17\,{\rm M}_{\odot}$ and $P_{\rm orb}=11.03$ hr; Papitto et al., 2013), and XSS J12270-4859 ($M_{\rm c,min}=0.15\,{\rm M}_{\odot}$ and $P_{\rm orb}=6.91$ hr; Roy et al., 2015; de Martino et al., 2015; Bassa et al., 2014).
@@
Figure 7: Orbital period versus companion mass for diverging binary tracks from our simulated grids that have a similar NS spin as the observed PSR J0952–0607 (1.41 ms) within a 10 % margin (dashed gray curves). The term $M_{\rm c}$ is the mass of the companion to the NS, and $P_{\rm orb}$ is the orbital period. All the tracks shown have $\beta=0.3$ and $f_{\rm pulsar}=0.5$. We further show the binary that best matches the sources in its orbital evolution in blue curve, which has initial companion mass of 2.0 ${\rm M}_{\odot}$, initial NS mass of 1.3 ${\rm M}_{\odot}$, and initial orbital period of 1.0 d.
@@ 3.3 Effect of pulsar wind irradiation
Figure 1 shows the evolutionary grids with an $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$ and varying pulsar wind efficiency ($f_{\rm pulsar}$) and we see the need for strong irradiation in order to explain all of the observed spiders. We also show the initial parameter space for spiders in Figure 4, their dynamical end, as well as the final orbital periods in the color bar, for $f_{\rm pulsar}=0.0$ and 0.5 and initial NS masses $1.3\,{\rm M}_{\odot}$. Many of the binaries shown in Figure 4 (around 50 to 60% in all simulated spiders) end up suffering a dynamical instability (shown by triangle symbols in the figure) divided into two categories; those that become unstable during the first RLO phase and those that become unstable toward the end of the second RLO phase. For these binaries in the second category, the NS still manages to accrete enough material to spin it up to milliseconds during the preceding case A RLO. Hence, these could be seen as spider binaries until the onset of instability, after which they would be seen as isolated spinning pulsars. Figure 5 shows the total mass-loss rates and the mass-transfer rates of the companion versus time of three types of interacting binaries (with pulsar wind irradiation) we have studied so far: case A , near-$P_{\rm bif}$, and case B binaries. The binaries have initial orbital properties as $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$ and $M^{i}_{\rm NS}=1.3\,{\rm M}_{\odot}$, and three values for $P^{i}_{\rm orb}=$1.0, 2.4, and 2.6 d. Going from low irradiation ($f_{\rm pulsar}=0.05$) to high ($f_{\rm pulsar}=0.5$), we see the mass-loss rates spike to very high values for case A and near-$P_{\rm bif}$ binaries, which could instigate an instability.
@@
We discussed the occurrence of dynamical instability toward the end of the second RLO phase in the previous section. The presence of irradiation would only increase the occurrence of this instability. Irradiation removes matter from the surface of the companion star, causing rapid stellar expansion. Increased $f_{\rm pulsar}$ results in a higher number of dynamical unstable systems (goes from 50% to 62% in simulated spiders when going from 0.0 to 0.5 in $f_{\rm pulsar}$). We include these dynamically unstable systems in our study since by the time the instability onset happens, the average MSP binary is already $\sim 10^{\color[rgb]{0,0,0}10}\rm\,yrs$ old and could have been observed during the preceding MSP phase. Hence, there is a reasonable chance to observe these binaries as BW binaries within a Hubble time ($1.4\times 10^{10}\rm\,yrs$). For this part of the parameter space the companion star would be destroyed due the instability. Therefore, it could provide an additional channel to the population of Galactic isolated MSPs (20% of the total Galactic MSP population) that is observed (Manchester et al., 2005a), with previous RLO spinning up the NS and subsequent merging destroying the companion. The idea that isolated MSPs could originate from BWs irradiating away their companions has been previously suggested both theoretically (van den Heuvel & van Paradijs, 1988) and through observations of MSPs with planets/asteroid belts (Wang et al., 2009; Shannon et al., 2013).
@@
Going from $f_{\rm pulsar}=0.0$ to $f_{\rm pulsar}=0.5$ in the converging systems (Figure 4), the final orbital periods widen significantly and this effect depends on the initial properties of binary. The widest orbits are seen with $M^{i}_{\rm c}=1.0\,{\rm M}_{\odot}$, where the final orbital periods are $\gtrsim 10$ d. For higher companion masses, the final orbital periods lie in the range of 0.1 to 1 d. Figure 6 shows the orbital evolution corresponding to three $M^{i}_{\rm c}$ (1.0, 1.5, and 2.5 ${\rm M}_{\odot}$) and $f_{\rm pulsar}=0.0$ to $f_{\rm pulsar}=0.5$, for various values of $P^{i}_{\rm orb}$. For case A RLO binaries ($P^{i}_{\rm orb}=1.0$ d; shown by black lines), we see increased orbital widening with $f_{\rm pulsar}=0.5$. The companions to the NS become fully convective around 0.2 ${\rm M}_{\odot}$, after which the pulsar wind is turned on. The orbit responds to the mass loss by expanding depending on $f_{\rm pulsar}$. Going from $M^{i}_{\rm c}=1.0\,{\rm M}_{\odot}$ to $1.5\,{\rm M}_{\odot}$ (for $f_{\rm pulsar}=0.5$) the expanding orbits cover a larger part of the observed BW space. This is due the fact that all of these binaries are undergoing RLO on a nuclear timescale. For a fixed NS mass and orbital period, more massive companions (or higher $q$) lead to more accretion onto the NS and a higher spin-down luminosity (and correspondingly high orbital expansion). However, with an $M^{i}_{c}=4.0\,{\rm M}_{\odot}$, the accreted matter decreases down to about $0.004\,{\rm M}_{\odot}$ as the binary suffers a dynamical instability soon after the first RLO-onset. This is significantly less than $1.0\,{\rm M}_{\odot}$ accreted with $M^{i}_{c}=3.0\,{\rm M}_{\odot}$. Hence, there is a limit to how high in companion mass we can go with a fixed initial NS mass and reproduce BWs. We can see this clearly by looking at Figure 4. The effect of irradiation on RLO can be seen in Figure 5 (top row) with increasing mass loss after the first RLO phase with increasing $f_{\rm pulsar}$.
@@
Next we look at the tracks for $P^{i}_{\rm orb}=2.4$ d in Figure 6, corresponding to the gray curves (left panel). For $M^{i}_{\rm c}=1.5\,{\rm M}_{\odot}$, the effect of $f_{\rm pulsar}=0.5$ is significant on the orbital evolution. The near-$P_{\rm bif}$ binary maintains a relatively wide orbit (around 0.2 d) crossing the RB regime. With increasing $f_{\rm pulsar}$, the duration of the mass-transfer phase is shortened as majority of mass loss is due to irradiation (see Figure 5). The radio pulsar wind will start as long as the two conditions described in Section 2.3.3 are met. Hence, strong irradiation disrupts the RLO phase as the companion detaches more easily from its Roche lobe due to higher mass loss. This does not significantly reduce the number of spiders binaries produced because most of the accretion happens early on in the evolution. In Figure 3, we see the NSs spin-up to milliseconds toward the start of RLO, for all the binaries presented. Going to $M^{i}_{\rm c}=2.5\,{\rm M}_{\odot}$ and $P^{i}_{\rm orb}=2.0$ d (gray curves in right panel), the near-$P_{\rm bif}$ tracks no longer pass through the RB region with the heavier companion (irrespective of $f_{\rm pulsar}$). For intermediate-mass companions ($\gtrsim 2.0\,{\rm M}_{\odot}$), the initial angular momentum loss by RLO dominates, as magnetic braking is not operating. For very short durations, the mass-transfer rate could even breach the Eddington limit, which results in any excess material being lost from the system taking away even more angular momentum. This continues until the star reaches around $1.0\,{\rm M}_{\odot}$ (and 1 d in Figure 6) and magnetic braking turns on. The following evolution is governed by magnetic braking losses and behaves similar to case A RLO, during its first RLO phase. The fast initial mass loss causes the initially $2.5\,{\rm M}_{\odot}$ star to become fully convective around $0.2\,{\rm M}_{\odot}$, after which it follows a UCXB-like evolution. Hence, initial companion masses $\gtrsim 2.0\,{\rm M}_{\odot}$ in near-$P_{\rm bif}$ binaries do not reproduce RBs (also shown in Figure 4). If a heavy NS is present ($\sim 2.0\,{\rm M}_{\odot}$), it could be argued that correspondingly heavier companions ($\gtrsim 2.0\,{\rm M}_{\odot}$) could form RBs. However, NSs with masses $1.8\,{\rm M}_{\odot}$ and above very easily cross the maximum NS mass threshold and very few binaries in those grids would be viable spiders (see Section 3.1).
@@
Figure 8: Final NS spins ($P^{f}_{\rm spin}$) versus final NS masses ($M^{f}_{\rm NS}$) in diverging binaries for $\beta=0.0$ (blue colored symbols), $\beta=0.3$ (orange colored symbols), and $\beta=0.7$ (green colored symbols). We group the binaries by the initial NS mass ($M^{i}_{\rm NS}$) corresponding to different symbols (see legend). We also compare to the observed spin and estimated pulsar mass of the huntsman source 1FGL J1417.7–440 (2.66 ms and $1.62^{+0.43}_{-0.17}\,{\rm M}_{\odot}$, shown by the gray dashed lines and gray shaded area for the uncertainties; Camilo et al., 2016; Swihart et al., 2018) and the estimated pulsar mass of the huntsman candidate 2FGL J0846.0+2820 ($1.96^{+0.41}_{-0.41}\,{\rm M}_{\odot}$, shown by the vertical pink dashed lines and pink shaded area for the uncertainties; Swihart et al., 2017) .
Figure 9: Allowed initial parameter space to form huntsmen spiders ($f_{\rm pulsar}=0.0)$ for an initial NS of mass $1.3\,{\rm M}_{\odot}$ (left) and $2.0\,{\rm M}_{\odot}$ (right). The term $M^{i}_{\rm c}$ is the initial mass of the companion to the NS, and $P^{i}_{\rm orb}$ is the initial orbital period. The color bar shows the final NS spins ($P^{f}_{\rm spin}$), and the gray stars denote the systems that formed diverging orbits but did not spin up to the MSP range. We show the end states of the tracks, if they encountered a dynamical instability (triangle symbols) or ended with one of the criteria for successful termination (star symbols). Converging tracks are shown by gray crosses.
Table 1: Observed and estimated properties of the two observed huntsmen spiders.
@@
 [^1] The term $M_{\rm c,min}$ is the minimum mass, calculated assuming a pulsar mass of $1.4\,{\rm M}_{\odot}$ and inclination of $90\degree$; $P_{\mathrm{orb}}$ is the orbital period; $P_{\rm spin}$ is the observed pulsar spin; $M_{\mathrm{NS}}$ is the estimated pulsar mass; $M_{\mathrm{c}}$ is the companion mass observed from Fermi-LAT; and $e$ is the observed eccentricity.